Hello.
I'm having a hard time getting this answer correct on homework
Solve the equation for a.
(a + 4)/b = (a - 4)/b + (b + 4)/a
a =
thanks!

Answer by stanbon(75887)

Solve the equation for a.
(a + 4)/b = (a - 4)/b + (b + 4)/a
----------------------
Multiply thru by ab to get:
a(a+4) = a(a-4) + b(b+4)
a^2 + 4a = a^2 - 4a + b^2 + 4b
Subtract a^2 from both sides
Add 4a to both sides
8a = b^2+4b
a = (b^2+4b)/8
===================================
Cheers,
Stan H.